OpenStudy (anonymous):

Find the equation, in slope-intercept form, for the line perpendicular to the given line and passing through the indicated point.

OpenStudy (anonymous):

Perpendicular to 2(y+x)-3(x-y)=-9 and passing through (2,5)

OpenStudy (anonymous):

Ok, first we need the slope of this given line, m1. The line perpendicular to the given line will have a slope m2, where m1m2=-1 expanding the above: 2y+2x-3x+3y=-9 5y=x-9 y=(1/5)-(9/5) Okay so m1=1/5 Thus (1/5)m2=-1 m2=-5 We need a line with slope=-5 through (2,5). y-5=-5(x-2) y-5=-5x+10 y=-5x+15

OpenStudy (anonymous):

Wow good job!

OpenStudy (anonymous):

ty :)

OpenStudy (anonymous):

Welcome :)